# Computing Perfect Bayesian Equilibria, with Application to Empirical Game-Theoretic Analysis

### v1

@@ Abstract.
Perfect Bayesian Equilibrium (PBE) is a refinement of the Nash equilibrium for imperfect-information extensive-form games (EFGs) that enforces consistency between the two components of a solution: agents’ strategy profile describing their decisions at information sets and the belief system quantifying their uncertainty over histories within an information set. We present a scalable approach for computing a PBE of an arbitrary two-player EFG. We adopt the definition of PBE enunciated by Bonanno in 2011 using a consistency concept based on the theory of belief revision due to Alchourrón, Gärdenfors and Makinson. Our algorithm for finding a PBE is an adaptation of Counterfactual Regret Minimization (CFR) that minimizes the expected regret at each information set given a belief system, while maintaning the necessary consistency criteria. We prove that our algorithm is correct for two-player zero-sum games and has a reasonable slowdown in time-complexity relative to classical CFR given the additional computation needed for refinement. We also experimentally demonstrate the competent performance of PBE-CFR in terms of equilibrium quality and running time on medium-to-large non-zero-sum EFGs. Finally, we investigate the effectiveness of using PBE for strategy exploration in empirical game-theoretic analysis. Specifically, we compute PBE as a meta-strategy solver (MSS) in a tree-exploiting variant of Policy Space Response Oracles (TE-PSRO). Our experiments show that PBE as an MSS leads to higher-quality empirical EFG models with complex imperfect information structures compared to MSSs based on an unrefined Nash equilibrium.
@@ 1. Introduction
Bonanno bonanno_plaus11; agm11 introduced a new consistency notion that covers both on- and off-equilibrium paths, based on the theory of belief revision due to Alchourrón, Gärdenfors and Makinson AGM85, that he termed AGM-consistency after the original authors. This notion requires the concept of plausibility orders over the nodes (representing histories or, equivalently, decision points) of the EFG tree based entirely on structural properties (edge incidence and information set membership). An assessment is AGM-consistent if it is possible to construct a plausibility order such that positive probabilities are assigned to nodes or edges of the EFG tree by the assessment if and only if they satisfy certain relationships under the order in question. agm11 reused the same term ‘perfect Bayesian equilibrium’ (PBE) for this refinement using AGM-consistency, which is still weaker than KW-consistency but easier to algorithmically verify, in principle—we will use this definition of PBE in this paper. Although it is known that every finite EFG admits at least one PBE agm11, the implementation and evaluation of robust, scalable algorithmic approaches towards the computation of a PBE for arbitrary dynamic games of imperfect information is an important practical question. This is the primary research question that motivates this work.
The framework of empirical game-theoretic analysis (EGTA) (wellman2025empirical) is highly relevant to our work. For multiagent scenarios that are too complicated for an analytic representation but admit procedural descriptions that can be queried (e.g., a simulator), EGTA offers a toolkit for using data collected from such queries to estimate a coarser model, called an empirical game; a key idea is to make this model amenable to off-the-shelf game-solving algorithms so that approximate insights about the underlying scenario can be derived from it. A popular and powerful iterative approach to EGTA called policy space response oracles (PSRO) psro17 uses an arbitrary game-solving algorithm as a module called the meta-strategy solver (MSS), which provides a principled basis for exploring the underlying strategy space to augment the model. Although the empirical game in EGTA has traditionally been maintained in the less expressive normal form, konicki22; konicki25 developed a tree-exploiting variant of EGTA (TE-EGTA) that represents the empirical game as an EFG tree. This approach makes use of refined solution concepts feasible for empirical games and potentially conducive to higher-quality game models, and the MSS for a tree-based model provides a natural use case for our algorithmic contributions.
@@ 1.1. Our Contributions
  • For two-player zero-sum games, we prove that the algorithm is correct by establishing a guarantee of convergence to an exactly sequentially rational solution and analyze its space and time complexity (Section 4).
@@ 1.2. Further Related Work
We provide a more detailed review of consistency concepts for NE refinements for EFGs in Apps. sections H.1 to H.3. A body of work exists on algorithms for computing (approximate) NE refinements (azhar05; turocy2010computing; panozzo14; pbe_auc_vinzenz23; graf_sequent_eq24), but the scalability and practicality of these algorithms has not been adequately established, to the best of our knowledge; please see App. H.4 for further details. There is a rich literature on extensions of the CFR approach including warm-start CFR brown_cfr14, CFR+ cfrplus14, CFR-D zun_subgame4, discounted CFR brown2019superhuman, linear CFR brown2019solving, deep CFR dcfrm_19, Monte Carlo CFR mccfr09, PCFR+ farina2021faster, and dynamic discounted CFR xu2024dynamic. One motivation for our work comes from the scalable, modular implementation of the generalized backward induction algorithm kaminski19 by konicki25 for computing an SPE of an imperfect-information EFG; they found in experiments that TE-PSRO with an SPE as MSS converges to a high-quality model faster than with an unrefined NE as MSS for diverse game classes. For an overview of EGTA techniques, including PSRO, we refer the reader to the survey by wellman2025empirical.
@@ 5.1. Experimental Setup
In PSRO, the model is updated in the following steps, illustrated in Figure 2). First, an arbitrary game-solving algorithm, called the meta-strategy solver (MSS) in this context, is applied to the current empirical game to obtain a solution called the target. Then, each agent’s best response (BR) to this target is approximated by a single-agent deep reinforcement learning (DRL) approach in an environment consistent with the simulator. Finally, agents’ strategy sets are augmented with the respective BRs, and the simulator is queried to obtain further data (payoff information, in particular) to complete the latest empirical game iterate.
Moreover, a game-solver termed EVAL, not necessarily the same as the MSS, gauges model quality and decides whether further refinement is necessary. konicki22 introduced the Tree-Exploiting PSRO (TE-PSRO) variant where the empirical game is in extensive form, though still a coarser version of a full description of the true game. konicki25 made methodological advances in this domain to improve the tradeoff between tractability and fidelity of the induced empirical game, and hence the scalability of TE-PSRO for imperfect-information games. They devised an abstraction framework in which each edge of the extensive-form empirical game represents a DRL-derived implicit policy executable in the simulator, allowing much of the underlying state and observation spaces to remain implicit in the model. They also employed a parameterized heuristic to control the growth of the empirical game tree by adding edges induced by the latest BR policies at select information sets only. For a fixed integer $M$, they first estimated the gains of playing BR policies rather than target policies at candidate information sets of the current model, constructed a softmax distribution over these information sets using those gains, and then sampled (up to) $M$ information sets from this distribution for adding edges to. We adopt this framework and call $M$ the growth parameter.
@@ 6. Discussion
This work was supported in part by a grant from the Effective Altruism Foundation and by the US National Science Foundation under CRII Award 2153184.
@@ Appendix F Bargain: Sequential Bargaining Game with Outside Offers
This game was adapted by konicki25 from a well-known alternating-offer bargaining game of incomplete information that has been used in a variety of AI studies, even featuring the use of large-language models (devault_dond15; lewis_dond17; strouse_dond21; li_dond23). They stripped away all natural-language elements from most previous implementations but added an external fallback option to augment the informational and strategic complexity.
@@ H.4. Other Related Work
More recently, pbe_auc_vinzenz23 introduced a best-response based algorithm for computing $\varepsilon$-perfect Bayesian equilibria in sequential auctions with incomplete information and combinatorial bidding; since auctions are a form of multi-stage game, this work demonstrated how to algorithmically compute an approximate PBE-MSG. graf_sequent_eq24 introduced an algorithm that symbolically solved imperfect-information games for sequential equilibria using a finite system of polynomial equations and inequalities; however, the algorithm is viable only for small games and not scalable.

### v2

@@ Abstract.
Perfect Bayesian Equilibrium (PBE) is a refinement of the Nash equilibrium for imperfect-information extensive-form games (EFGs) that enforces consistency between the two components of a solution: agents’ strategy profile describing their decisions at information sets and the belief system quantifying their uncertainty over histories within an information set. We present a scalable approach for computing a PBE of an arbitrary two-player EFG. We adopt the definition of PBE enunciated by Bonanno in 2011 using a consistency concept based on the theory of belief revision due to Alchourrón, Gärdenfors, and Makinson. Our algorithm for finding a PBE is an adaptation of Counterfactual Regret Minimization (CFR) that minimizes the expected regret at each information set given a belief system, while maintaining the necessary consistency criteria. We prove that our algorithm is correct for two-player zero-sum games and has a reasonable slowdown in time-complexity relative to classical CFR given the additional computation needed for refinement. We also experimentally demonstrate the competent performance of PBE-CFR in terms of equilibrium quality and running time on medium-to-large non-zero-sum EFGs. Finally, we investigate the effectiveness of using PBE for strategy exploration in empirical game-theoretic analysis. Specifically, we compute PBE as a meta-strategy solver (MSS) in a tree-exploiting variant of Policy Space Response Oracles (TE-PSRO). Our experiments show that PBE as an MSS leads to higher-quality empirical EFG models with complex imperfect information structures compared to MSSs based on an unrefined Nash equilibrium.
@@ 1. Introduction
Bonanno bonanno_plaus11; agm11 introduced a new consistency notion that covers both on- and off-equilibrium paths, based on the theory of belief revision due to Alchourrón, Gärdenfors, and Makinson AGM85, which he termed AGM-consistency after the original authors. This notion requires the concept of plausibility orders over the nodes (representing histories or, equivalently, decision points) of the EFG tree based entirely on structural properties (edge incidence and information set membership). An assessment is AGM-consistent if it is possible to construct a plausibility order such that positive probabilities are assigned to nodes or edges of the EFG tree by the assessment if and only if they satisfy certain relationships under the order in question. agm11 reused the same term ‘perfect Bayesian equilibrium’ (PBE) for this refinement using AGM-consistency, which is still weaker than KW-consistency but easier to algorithmically verify in principle—we will use this definition of PBE in this paper. Although it is known that every finite EFG admits at least one PBE agm11, the implementation and evaluation of robust, scalable algorithmic approaches towards the computation of a PBE for arbitrary dynamic games of imperfect information is an important practical question. This is the primary research question that motivates this work.
The framework of empirical game-theoretic analysis (EGTA) (wellman2025empirical) is highly relevant to our work. For multiagent scenarios that are too complicated for an analytic representation but admit procedural descriptions that can be queried (e.g., a simulator), EGTA offers a toolkit for using data collected from such queries to estimate a coarser model, called an empirical game; a key idea is to make this model amenable to off-the-shelf game-solving algorithms so that approximate insights about the underlying scenario can be derived from it. A popular and powerful iterative approach to EGTA called policy space response oracles (PSRO) bighashdel_psro24; psro17 uses an arbitrary game-solving algorithm as a module called the meta-strategy solver (MSS), which provides a principled basis for exploring the underlying strategy space to augment the model. Whereas the empirical game in EGTA has commonly been maintained in the less expressive normal form, we developed a tree-exploiting variant of EGTA (TE-EGTA) in prior work (konicki22; konicki25) that represents the empirical game as an EFG tree. This approach makes use of refined solution concepts feasible for empirical games and potentially conducive to higher-quality game models, and the MSS for a tree-based model provides a natural use case for our algorithmic contributions in this paper.
@@ 1.1. Our Contributions
  • For two-player zero-sum games, we prove that the algorithm is correct by establishing a guarantee of convergence to an exactly sequentially rational solution, and analyze its space and time complexity (Section 4).
@@ 1.2. Further Related Work
We provide a more detailed review of consistency concepts for NE refinements for EFGs in Apps. sections H.1 to H.3. wellman2025empirical provides an overview of EGTA techniques including PSRO. A body of work exists on algorithms for computing (approximate) NE refinements (azhar05; turocy2010computing; panozzo14; Thoma25; graf_sequent_eq24), but the scalability and practicality of these algorithms has not been adequately established, to the best of our knowledge; please see App. H.4 for further details. There is a rich literature on extensions of the CFR approach including warm-start CFR brown_cfr14, CFR+ cfrplus14, CFR-D zun_subgame4, discounted CFR brown2019superhuman, linear CFR brown2019solving, deep CFR dcfrm_19, Monte Carlo CFR mccfr09, PCFR+ farina2021faster, and dynamic discounted CFR xu2024dynamic. In prior work, we developed a scalable, modular implementation konicki25 of the generalized backward induction algorithm kaminski19 for computing an SPE of an imperfect-information EFG, and found in experiments that TE-PSRO with an SPE as MSS converges to a high-quality model faster than with an unrefined NE as MSS for diverse game classes. A similar treatment of PBE is a natural next step.
@@ 5.1. Experimental Setup
In PSRO, the model is updated in the following steps, illustrated in Figure 2). First, an arbitrary game-solving algorithm, called the meta-strategy solver (MSS) in this context, is applied to the current empirical game to obtain a solution called the target. Then, each agent’s best response (BR) to this target is approximated by a single-agent deep reinforcement learning (DRL) approach in an environment consistent with the simulator. Finally, agents’ strategy sets are augmented with the respective BRs, and the simulator is queried to obtain further data (payoff information, in particular) to complete the latest empirical game iterate. Moreover, a game-solver termed EVAL, not necessarily the same as the MSS, gauges model quality and decides whether further refinement is necessary.
In prior work konicki22, we introduced the Tree-Exploiting PSRO (TE-PSRO) variant where the empirical game is in extensive form, though still a coarser version of a full description of the true game. We followed up with methodological advances to improve the tradeoff between tractability and fidelity of the induced empirical game, and hence the scalability of TE-PSRO for imperfect-information games konicki25. We devised an abstraction framework in which each edge of the extensive-form empirical game represents a DRL-derived implicit policy executable in the simulator, allowing much of the underlying state and observation spaces to remain implicit in the model. We also employed a parameterized heuristic to control the growth of the empirical game tree by adding edges induced by the latest BR policies at select information sets only. For a fixed integer $M$, we first estimated the gains of playing BR policies rather than target policies at candidate information sets of the current model, constructed a softmax distribution over these information sets using those gains, and then sampled (up to) $M$ information sets from this distribution for adding edges to. We adopt this framework in this paper too and call $M$ the growth parameter.
@@ 6. Discussion
This work was supported in part by the US National Science Foundation under CRII Award 2153184.
@@ Appendix F Bargain: Sequential Bargaining Game with Outside Offers
We adapted this game from a well-known alternating-offer bargaining game of incomplete information that has been used in a variety of AI studies, even featuring the use of large-language models (devault_dond15; lewis_dond17; strouse_dond21; li25lmmghlbw). The adapted version strips away all natural-language elements from the original, but adds an external fallback option to augment the informational and strategic complexity.
@@ H.4. Other Related Work
More recently, Thoma25 introduced a best-response based algorithm for computing $\varepsilon$-perfect Bayesian equilibria in sequential auctions with incomplete information and combinatorial bidding; since auctions are a form of multi-stage game, this work demonstrated how to algorithmically compute an approximate PBE-MSG. graf_sequent_eq24 introduced an algorithm that symbolically solved imperfect-information games for sequential equilibria using a finite system of polynomial equations and inequalities; however, the algorithm is viable only for small games and not scalable.